# A Complexity-Based Theory of Compositionality

### v1

@@ 4.3 Natural languages
To investigate the compositionalities of natural language systems using our definition, we first collected a dataset of English sentences describing natural images (COCO,, 2024), which we then translated into French, Spanish, German, and Japanese using a large open source model (Costa-jussà et al.,, 2022). To obtain proxies of “meanings” $Z$ for these sentences, we encoded them using a multilingual sentence embedding model that outputs a dense fixed-size vector (Reimers and Gurevych,, 2020). More experimental details as well as limitations of this approach can be found in Appendix H. Using these datasets of sentence/representation pairs, we measured the compositionalities of each natural language system $C^{L}(Z)$ using the same prequential coding approach as in Section 4.2.
@@
Our results are shown in Figure 4. We find that the prequential code lengths of all languages are highly similar, indicating that they have semantics $f$ of roughly equal complexity (Figure 4a). Assuming that these natural languages are all equally expressive in their abilities to express ideas and identify referents (i.e., equal $K(Z)$; a common assumption in linguistics), their compositionalities as measured by our definition $C^{L}(Z)$ are roughly equivalent, with Japanese having slightly higher relative compositionality (Figure 4b). Using topological similarity as an alternative definition of compositionality gives counter-intuitive results that contradict our own: most languages have a near-zero topological similarity, except for Japanese which is a strong outlier with a topological similarity of $-0.2$ (Figure 4c).
@@ 5 Conclusion
In virtue of being quantitatively precise, representational compositionality can be used to investigate compositionality in real-world systems. We demonstrated this in the case of emergent and natural language representations, but in a limited way that only considered language systems where the sentences describing a representation are externally defined. We note that this quantity can readily be applied to score tokenization schemes and may help guide improvements in tokenizer design, but even more intriguing is the possibility to operationalize our proposed measure $C^{L}$ to regularize and otherwise bias next generation tokenizers. We leave the exploration of this topic to future work. More generally however, measuring the compositionality of representations without a predefined mapping to sentences requires the development of additional machine learning tools, who’s overall architecture we sketch out in Appendix B. The development of such tools is an important direction for future work, as it will allow us to investigate the compositionality of representations that emerge from different learning objectives, neural architectures, inductive biases, and brain regions. Further, these insights could potentially be used to validate or reject hypotheses about the role of compositionality in human cognition, such as the Language of Thought hypothesis (Fodor,, 1975).
Representational compositionality can also play an important role in the design and validation of machine learning models with principled inductive biases for compositionality. Namely, in addition to supporting a given task, a compositional representation must be easily describable as a simple function of constituent parts. There are both direct and indirect ways to achieve this that are grounded in our definition, both of which we intend to pursue in future work.
@@ Appendix A Background on Kolmogorov complexity
$K(X|p)\leq K(X|p_{\theta})\leq-\sum_{i=1}^{n}\log_{2}p_{\theta}(x_{i}),$ (8)
where we have equality on the LHS when $p_{\theta}=p$ and equality on the RHS when the cost of improving $p_{\theta}$ (in bits of written code) would be greater than the benefits from more accurate modeling. In practice, if $p_{\theta}$ is close to $p$, we can say that $K(X|p_{\theta})\approx-\sum_{i=1}^{n}\log_{2}p_{\theta}(x_{i})$.
This insight is significant. Notice that $-\sum_{i=1}^{n}\log_{2}p_{\theta}(x_{i})$ is the negative log-likelihood of the data under the model, which is a common loss function used in ML. This tells us that models with lower error better compress their data, and directly relates Kolmogorov complexity to optimization in ML. However, what if we do not have a model? What is the Kolmogorov complexity of the data itself? Intuitively, if the dataset is sufficiently large, the optimal method for encoding it should be to first specify a model and then encode the data using that model as in Equation 8. Specifically, using identities in Fortnow, (2000), we have:
@@
This encoding scheme on the RHS is referred to as a 2-part code (Grünwald,, 2007). For large datasets, we have equality when the model’s description length and error are jointly minimized, which occurs when the model $p_{\theta}(x)$ is equivalent to the true distribution $p(x)$:
@@ Appendix B Compressing a representation using discrete auto-encoders
$=\min_{p_{w},W,f}K(p_{w})-\sum_{n=1}^{N}\log p_{w}(w_{n})+K(f)-\sum_{n=1}^{N}\log p(z_{n};f(w_{n})).$
While this is an intractable search problem, it can be turned into an easier optimization problem using modern deep learning tools. In particular, we can minimize at least some of the terms in Equation 1 by fitting a discrete auto-encoder to $Z$ using a learned prior in the latent $W$-space, as illustrated in Figure B.1. This auto-encoder consists of an encoder $w=e(z)$ that maps the representation to a discrete latent space of sentences, a latent prior $p_{w}(w)$, and a decoder $p(z;f(w))$ that outputs the sufficient statistics of a Gaussian distribution in order to evaluate the likelihood of the original representation. In practice, the latent prior $p_{w}(w)$ can be parameterized using an auto-regressive model such as a causal Transformer, which tends to work well on language data. We can then train this discrete auto-encoder using the following loss function:
$\mathcal{L}(Z;e,p_{w},f)=\sum_{z\in Z}-\log p_{w}(e(z))-\log p(z;f(e(z))).$ (12)
@@
Regardless of which method is used, the complexities of the final trained models can be estimated using a method called prequential coding (Blier and Ollivier,, 2018), which we describe in Appendix E. Thus, we are able to estimate all of the constituent complexity terms of $K(Z)$ in Equation 1. The main challenge in this overall approach then becomes how to successfully train a discrete auto-encoder with a prior in latent space, in a way that is both stable and scalable.
@@ Appendix D Examples of compositional representations
To supplement and clarify the arguments in Section 3, it is easiest to gain further intuition for our definition of compositionality through concrete examples of different hypothetical representations. For each, we have strong intuitions about whether or not the representation is compositional, and we will see that our definition agrees with—and indeed extends—these intuitions.
@@
Example 2, $\downarrow C(Z)$: $Z$ is a smooth continuous function
@@
Most of the discussion thus far has focused on the denominator of $C(Z)$ in Definition 2. However, a representation can also lack compositionality if the complexity of the numerator, $K(Z)$, is low. If $Z$ were very low—say it were a constant, for instance—then it could be modeled using a simple $f$ that achieves low error $K(Z|W,f)$. However, we would certainly not be tempted say that the representation is compositional. In fact, it would be best compressed using a single word and an $f$ that outputs a constant, rather than using complex sentences and simple compositional rules. Compositionality must therefore also increase with the expressivity of the representation, which is captured by the numerator $K(Z)$.
@@
As already explained in Section 2, a modular $f$ is simpler to describe and thus implies higher compositionality. This accounts for intuitions about the relationship between compositionality and models that exhibit structural modularity (Lepori et al.,, 2023; Goyal and Bengio,, 2022).
@@
The connection between equivariance and compositionality is perhaps less obvious (Gordon et al.,, 2020), but it is a natural and intuitive consequence of our definition. Equivariance (and invariance) is a source of structure that decreases the complexity of a function (Immer et al.,, 2022; Wilk et al.,, 2018; van der Ouderaa and van der Wilk,, 2022). For instance, convolutional layers have local connectivity and reuse weights across spatial locations, which both reduces their description length and makes them equivariant to spatial translations. We can also consider linear equivariance as a special case that is easy to illustrate. If $f$ is linearly equivariant to a particular operation $g$ in sentence-space, it means that $f(g(w))=f(w)+v_{g}$, where $v_{g}$ is a constant vector that corresponds to the equivariant change in the representation output by $f$. The difference in the function’s behaviour for two different inputs, $w$ and $g(w)$, can therefore be compactly described by a single vector, whereas in the general non-equivariant case the change in the function’s behaviour can be arbitrarily complex. In an extreme case, if $f$ can be completely described by a set of linear equivariances, then each $w$ corresponds to a set of $g_{i}$’s applied to a constant “default” sentence, and $f$ merely needs to encode a single vector for each of these $g_{i}$’s then sum those that apply to a particular input. The resulting function is very similar to the one described in Example 4, where $f$ applied a simple operation to a sequence of word embeddings in a sentence (in this case vector addition). The function also bears similarities to the one described in Example 5 if we view the equivariances as modules. Similar arguments can be made for non-linear equivariance, where the complexity $K(f)$ would still be reduced, but to a lesser extent. In general, the more equivariances a function has and the simpler those equivariances are, the lower the complexity $K(f)$ and the higher the compositionality $C(Z)$.
Appendix E Prequential coding

### v3

@@ 4.3 Natural languages
To investigate the compositionalities of natural language systems using our definition, we first collected a dataset of English sentences describing natural images (COCO,, 2024), which we then translated into French, Spanish, German, and Japanese using a large open source model (Costa-jussà et al.,, 2022). To obtain proxies of “meanings” $Z$ for these sentences, we encoded them using a multilingual sentence embedding model that outputs a dense fixed-size vector (Reimers and Gurevych,, 2020). More experimental details as well as limitations of this approach can be found in Appendix J. Using these datasets of sentence/representation pairs, we measured the compositionalities of each natural language system $C^{L}(Z)$ using the same prequential coding approach as in Section 4.2.
@@
Our results are shown in Figure 4. We find that the prequential code lengths of all languages are highly similar, indicating that they have semantics $f$ of roughly equal complexity (Figure 4a). Assuming that these natural languages are all equally expressive in their abilities to express ideas and identify referents (i.e., equal $K(Z)$; a common assumption in linguistics), their compositionalities as measured by our definition $C^{L}(Z)$ are roughly equivalent, with Japanese having slightly higher relative compositionality (Figure 4b). Using topological similarity as an alternative definition of compositionality gives counter-intuitive results that contradict our own: most languages have a near-zero topological similarity, except for Japanese which is a strong outlier with a topological similarity of $-0.2$ despite having the lowest complexity mapping $K(Z|W)$ of all languages considered (Figure 4c).
@@ 5 Conclusion
In virtue of being quantitatively precise, representational compositionality can be used to investigate compositionality in real-world systems. We demonstrated this in the case of emergent and natural language representations where the sentences describing a representation are externally defined. We note that in future work this quantity can readily be applied to score tokenization schemes and may help guide improvements in tokenizer design, but even more intriguing is the possibility to operationalize our proposed measure $C^{L}$ to regularize and otherwise bias next generation tokenizers. We leave the exploration of this topic to future work.
More generally however, measuring the compositionality of representations without a predefined mapping to sentences requires the development of additional machine learning tools, whose overall architecture we sketch out in Appendix B. The development of such tools is an important direction for future work, as it will allow us to investigate the compositionality of representations that emerge from different learning objectives, neural architectures, inductive biases, and brain regions. In turn, we will be able to see how representational compositionality empirically relates to other topics in ML such as compositional generalization, multi-task generalization, and latent space generative models—we give some hypotheses and ideas for future work along these lines in Appendix E. In particular, representational compositionality has the potential to explain the success of varied methods because it defines compositionality through compression, which abstracts across the architecture, learning details, and particular representational format of a model. Representational compositionality can therefore be used to validate or reject diverse hypotheses about compositionality, such as the Language of Thought hypothesis (Fodor,, 1975).
Representational compositionality can also play an important role in the design and validation of machine learning models with principled inductive biases for compositionality. Namely, in addition to supporting a given task, a compositional representation must be easily describable as a simple function of constituent parts. There are both direct and indirect ways to achieve this that are grounded in our definition, and we describe some approaches in Appendix F that we intend to pursue in future work.
@@ Appendix A Background on Kolmogorov complexity
$K(X|p_{\theta})\leq-\sum_{i=1}^{n}\log_{2}p_{\theta}(x_{i}),$ (8)
where we have equality when $p_{\theta}=p$. This insight is significant. Notice that $-\sum_{i=1}^{n}\log_{2}p_{\theta}(x_{i})$ is the negative log-likelihood of the data under the model, which is a common loss function used in ML. This tells us that models with lower error better compress their data, and directly relates Kolmogorov complexity to optimization in ML. However, what if we do not have a model? What is the Kolmogorov complexity of the data itself? Intuitively, if the dataset is sufficiently large, the optimal method for encoding it should be to first specify a model and then encode the data using that model as in Equation 8. Specifically, using identities in Fortnow, (2000), we have:
@@
This encoding scheme on the RHS is referred to as a 2-part code (Grünwald,, 2007). We have equality when the model’s description length and error are jointly minimized, which occurs when the model $p_{\theta}(x)$ is equivalent to the true distribution $p(x)$:
@@ Appendix B Compressing a representation using discrete auto-encoders
$=\min_{p_{w},W,f}K(p_{w})-\sum_{n=1}^{N}\log p_{w}(w_{n})+K(f)-\sum_{n=1}^{N}\log\mathcal{N}(z_{n};f(w_{n})).$
While this is an intractable search problem, it can be turned into an easier optimization problem using modern deep learning tools. In particular, we can minimize at least some of the terms in Equation 1 by fitting a discrete auto-encoder to $Z$ using a learned prior in the latent $W$-space, as illustrated in Figure B.1. This auto-encoder consists of an encoder $w=e(z)$ that maps the representation to a discrete latent space of sentences, a latent prior $p_{w}(w)$, and a decoder $\mathcal{N}(z;f(w))$ that outputs the sufficient statistics of a Gaussian distribution in order to evaluate the likelihood of the original representation. In practice, the latent prior $p_{w}(w)$ can be parameterized using an auto-regressive model such as a causal Transformer, which tends to work well on language data. We can then train this discrete auto-encoder using the following loss function:
$\mathcal{L}(Z;e,p_{w},f)=\sum_{z\in Z}-\log p_{w}(e(z))-\log\mathcal{N}(z;f(e(z))).$ (12)
@@
Regardless of which method is used, the complexities of the final trained models can be estimated using a method called prequential coding (Blier and Ollivier,, 2018), which we describe in Appendix G. Thus, we are able to estimate all of the constituent complexity terms of $K(Z)$ in Equation 1. The main challenge in this overall approach then becomes how to successfully train a discrete auto-encoder with a prior in latent space, in a way that is both stable and scalable.
@@
Computational complexity
If the discrete auto-encoder described in this section can be trained successfully, then estimating representational compositionality is tractable, despite being defined theoretically in terms of Kolmogorov complexities. Fitting the auto-encoder itself is tractable using modern machine learning hardware. Then, to estimate $K(p_{w})$ and $K(f)$ we must use prequential coding (see Appendix G), which amounts to fitting a neural network at varying dataset sizes. While fitting a neural network $N$ times (where $N$ is the dataset size) is inefficient, it is nonetheless tractable, and can be approximated efficiently by chunking the data into coarser sizes as we did in our experiments. There are also methods for computing prequential coding online rather than retraining the model from scratch each iteration (Bornschein et al.,, 2022).
@@ Appendix D Examples of compositional representations
To supplement and clarify the arguments in Section 3, it is easiest to gain further intuition for our definition of compositionality through concrete examples of different hypothetical representations. For each, we have strong intuitions about whether or not the representation is compositional, and we will see that our definition agrees with—and indeed extends—these intuitions. We illustrate these examples in Figure D.1.
Figure D.1: Examples of different representations and their compositionalities according to $C(Z)$. Example 1. A representation whose clusters lack any structure has semantics $f$ that map $w\rightarrow z$ arbitrarily using a lookup table, resulting in high $K(f)$ and low $C(Z)$. Example 2. A representation that is smooth and continuous cannot be compressed as a function of discrete parts without incurring significant prediction error, resulting in high $K(Z|W,f)$ and low $C(Z)$. Example 3. A representation that cannot express many different things (thoughts, visual scenes, ideas, etc.), such as one that is sampled from a unimodal distribution, has low $K(Z)$ and low $C(Z)$. Example 4. A representation that can be described by assigning word embeddings which are then processed using a simple operation (e.g., concatenation, as in disentanglement) has low $K(f)$ and high $C(Z)$. Example 5. A representation whose semantics can be compressed using a small number of simple and reusable modules has low $K(f)$ and high $C(Z)$. Example 6. A representation whose semantics have a large number of symmetries, or equivariances, has low $K(f)$ and high $C(Z)$.
@@
Example 2, $\downarrow C(Z)$: $Z$ is smooth and continuous
@@
Most of the discussion thus far has focused on the denominator of $C(Z)$ in Definition 2. However, a representation can also lack compositionality if the complexity of the numerator, $K(Z)$, is low. If $Z$ were very low—say it were a constant, for instance—then it could be modeled using a simple $f$ that achieves low error $K(Z|W,f)$. However, we would certainly not be tempted say that the representation is compositional. In fact, it would be best compressed using a single word and an $f$ that outputs a constant, rather than using complex sentences and simple compositional rules. Compositionality must therefore also increase with the expressivity of the representation, which is captured by the numerator $K(Z)$ in our definition. In cognitive science, where the scientific notion of compositionality has its origins, expressivity is considered an essential component of compositionality; Chomsky, (1956) famously argued that natural language as a compositional system derives its power because it gives us “infinite use of finite means”, or in the language of our definition high expressivity as a simple function of parts.
@@
As already explained in Section 2, a modular $f$ is simpler to describe and thus implies higher compositionality. To see why modular functions are more compressible, consider a paradigmatic case: computer programs. When a computer program is written in such a way that it can be refactored into a small number of functions and classes that are reused several times, the total length of the program decreases substantially. Programs that are not written with modularity in mind tend to be much longer and complex. Modular functions therefore tend to have far lower complexity because the modules only need to be defined once, but can then be reused many times inside the function. In ML, modularity is leveraged in a similar fashion. For instance, Goyal et al., (2021) introduces an architecture that consists of $N$ DNNs as well as a learned attention-based routing mechanism for how they communicate. Crucially, these modules are leveraged by the routing mechanism in a context-dependent way, and each module can be reused many times to process each individual input. This means that while the entire model is simple (small number of modules and simple routing mechanism), it is nevertheless highly expressive due to the combinatorial way in which modules can be composed. Our definition explains how this expressivity and compression endowed by modular functions formally relates to compositionality (Lepori et al.,, 2023; Goyal and Bengio,, 2022).
@@
The connection between equivariance and compositionality is perhaps less obvious (Gordon et al.,, 2020), but it is a natural and intuitive consequence of our definition. Equivariances (and invariances) are symmetries—sources of structure that decreases the complexity of a function (Immer et al.,, 2022; Wilk et al.,, 2018; van der Ouderaa and van der Wilk,, 2022). For instance, convolutional layers have local connectivity and reuse weights across spatial locations, which both reduces their description length and makes them equivariant to spatial translations. We can also consider linear equivariance as a special case that is easy to illustrate. If $f$ is linearly equivariant to a particular operation $g$ in sentence-space, it means that $f(g(w))=f(w)+v_{g}$, where $v_{g}$ is a constant vector that corresponds to the equivariant change in the representation output by $f$. The difference in the function’s behaviour for two different inputs, $w$ and $g(w)$, can therefore be compactly described by a single vector, whereas in the general non-equivariant case the change in the function’s behaviour can be arbitrarily complex. In an extreme case, if $f$ can be completely described by a set of linear equivariances, then each $w$ corresponds to a set of $g_{i}$’s applied to a constant “default” sentence, and $f$ merely needs to encode a single vector for each of these $g_{i}$’s then sum those that apply to a particular input. The resulting function is very similar to the one described in Example 4, where $f$ applied a simple operation to a sequence of word embeddings in a sentence (in this case vector addition). The function also bears similarities to the one described in Example 5 if we view the equivariances as modules. Similar arguments can be made for non-linear equivariance, where the complexity $K(f)$ would still be reduced, but to a lesser extent. In general, the more equivariances a function has and the simpler those equivariances are, the lower the complexity $K(f)$ and the higher the compositionality $C(Z)$.
Appendix E Relations between representational compositionality and other ML topics
Compositional generalization
One of the benefits of compositional representations is that they enable better compositional generalization (Lake and Baroni,, 2018). If a model is compositional with respect to a set of features in its training data, it need not observe all possible combinations of those features in order to generalize to novel ones (Schug et al.,, 2024; Wiedemer et al.,, 2024, 2023; Bahdanau et al.,, 2019; Mittal et al.,, 2021; Hupkes et al.,, 2020; Jarvis et al.,, 2024; Lippl and Stachenfeld,, 2024; Lachapelle et al.,, 2024). For instance, if an image classifier’s representation is compositional with respect to foreground objects and background scenes, then it should be able to correctly classify an image of “a cow on a beach” at inference time after having only observed cows and beaches separately at training time.
In certain cases, compositionality is defined in terms of a model’s ability to compositionally generalize compositionally (e.g., Jarvis et al.,, 2024; Wiedemer et al.,, 2024, 2023; Lippl and Stachenfeld,, 2024). However, while such definitions of compositionality can often provide theoretical guarantees on generalization, they also place strong assumptions on either the representation, the downstream model, or both. For instance, Wiedemer et al., (2023) assumes that the representation is perfectly disentanglement with respect to some underlying task constituents. Similarly, Lachapelle et al., (2024) assumes disentanglement and that the downstream function using the representation is additive with respect to the the disentangled factors, and Lippl and Stachenfeld, (2024) assumes disentanglement and “conjunction-wise additivity”. Wiedemer et al., (2024) takes from the object-centric learning literature and defines a compositional representation as one that is structured into distinct “slots” (Locatello et al.,, 2020), and then requires that the downstream model using these slots is additive.
In contrast, our definition of representational compositionality is far more general: it defines compositionality in terms of compression, which abstracts across the architecture producing and using the representation, learning details, data requirements, and particular representational format. For instance, disentangled and slot-wise representations are particular cases of representational compositionality in terms of their simple semantics $K(f)$ (see Appendix D), but these are rigid assumptions to build into a model that might negatively impact performance. In contrast, representational compositionality has the potential to explain the success of more varied and flexible methods in terms of compositional generalization, such as loss regularizers or simply scaling dataset and model size.
As a consequence of its generality, it may be difficult to formally characterize the relationship between representational compositionality and compositional generalization with theoretical guarantees, and we did not attempt to do so in this paper. Nevertheless we hypothesize that representations with high $C(Z)$ should enable better compositional generalization. This is because the representation of constituent parts is systematic: the semantics mapping constituent parts to the representation is a simple function that will generalize better to novel part combinations (i.e., it will assign them a meaningful rather than arbitrary representation, which downstream functions should be able to leverage). One of our central goals for future work is to test this hypothesis empirically, where we measure the compositionalities of many model representations using our definition and then correlate this score with the models’ compositional generalization abilities.
Generative models in latent space
In addition to compositional generalization, representational compositionality also relates to generative models that sample in latent space. In particular, once a compositional representation is learned, efficient and generalizable generative models can be constructed by sampling in the space of discrete sentences, rather than in the high-dimensional continuous latent space directly. This is because the semantics function $f$ of a representation with high $C(Z)$ is simple, and can generalize to novel sentences that the generative model might produce. Empirically, modeling and sampling from discrete distributions is often easier and more effective, especially for complex multi-model distributions (Razavi et al.,, 2019).
To give a concrete example, imagine that a vision model has been pretrained on some task like object classification and produces latent representations with high $C(Z)$. Using this representation, we can train a generative model of the form $z\sim p_{w}(w)\mathcal{N}(z;f(w))$ described in Section 2, and then generate novel samples for downstream visual reasoning tasks directly in the abstract latent space, rather than in the low-level image space. This is similar to thought and reasoning involved in human cognition, which are generative processes believed to exhibit a discrete language-like structure (Fodor,, 1975; Dehaene et al.,, 2022; Lake et al.,, 2017; Bengio,, 2017; Goyal and Bengio,, 2022).
Appendix F Inductive biases for representational compositionality
In virtue of being formally precise and quantitative, representational compositionality can inspire the design of novel inductive biases for compositional representations in ML models. In this section, we outline two approaches that we believe have promise: one that directly optimizes for $C(Z)$, and another that indirectly attempts to increase it through task and data constraints. In addition, $C(Z)$ can be used to validate existing inductive biases for compositionality (e.g., architectures for object-centric representations Locatello et al.,, 2020).
Regularizing $K(Z|W)$
The most direct way to learn representations with high $C(Z)$ is to regularize the denominator $K(Z|W)$ so that the representations become more verbalizable, as suggested in Bengio, (2017) and Goyal and Bengio, (2022). Definition 2 says that compositional representations are (a) expressive and (b) easily described using sequences of discrete symbols—in other words, that they are verbalizable like human thoughts that can largely be conveyed in natural language. Expressivity can be obtained simply by training on a sufficiently complex task; for example, representations for image classification need to be expressive so that they can discriminate different objects. Task pressure alone, however, does not guarantee that the representation will be verbalizable. This second desiderata can be achieved, however, through a prior that regularizes the model’s loss function.
Say that some model $g_{\theta}$ produces a representation $Z=g_{\theta}(X)$ of inputs $X$. Verbalization corresponds to minimizing the denominator in Definition 2: $K(Z|W)=K(f)+K(Z|W,f)$. Crucially, $W$ and $f$ here are obtained from the shortest program that outputs $Z$ as described in Section 2, which can be approximated by optimizing a discrete auto-encoder who’s training scheme is sketched out in Appendix B. To make the dependence of $W$ and $f$ on $Z$ more explicit here, we will use the superscripts $W^{Z}$ and $f^{Z}$. If we wish to increase verbalizability (and therefore compositionality), we therefore need to perform some update update $\theta\rightarrow\theta^{\prime}$ such that:
$K(f^{Z^{\prime}})+K(Z^{\prime}|W^{Z^{\prime}},f^{Z^{\prime}})<K(f^{Z})+K(Z|W^{Z},f^{Z}),$ (13)
where $Z^{\prime}=g_{\theta^{\prime}}(X)$. One option for accomplishing this is by backpropagating the reconstruction error of the discrete auto-encoder, $K(Z|W^{Z},f^{Z})$. This approach assumes that the semantics before and after the update are unchanged (i.e., $f^{Z^{\prime}}=f^{Z}$), so that the only thing that needs to be considered is the auto-encoding reconstruction error $K(Z|W^{Z},f^{Z})\rightarrow K(Z^{\prime}|W^{Z^{\prime}},f^{Z})$. While this assumption will be violated in practice, it may hold approximately such that regularizing reconstruction error alone is sufficient to increase compositionality.
In sum, the approach described here consists of training a DNN $g_{\theta}(X)$ on some task as usual, but with an additional loss: a discrete auto-encoder is fit to a layer in the model which we want to be more compositional, and the $\theta$ is regularized to minimize the loss of this discrete auto-encoder. As a result, in addition to subserving task demands, the representation is optimized to be more compressible as a function of constituent discrete parts (i.e., it is verbalizable).
Multi-task training
A common observation in deep learning is that the model representations after training tend to be surprisingly simple despite the significant number of parameters in the network (Blier and Ollivier,, 2018), as evidenced by their strong iid generalization abilities. However, absent additional constraints (e.g., Lachapelle et al.,, 2024), these same representations do not enable compositional out-of-distribution generalization, suggesting that they lack sufficient compositional structure. One hypothesis is that while the simplest representation used to solve a single task may not be compositional, the simplest representation used to solve many related tasks might be. An analogy can be made to computer programs. When a program is written for a single narrow purpose, writing it in a compositional manner that reuses shared functions and classes might in fact result in bloat that increases the total program length. However, if these same functions and classes constitute a useful library that can be leveraged to write other programs as well, significant compression might be possible because the library is shared across all programs.
In the terminology of $C(Z)$, learning the simplest representation that subserves many different related tasks might result in low $K(Z|W)$ and high compositionality because the semantics $f$ are shared across these tasks and therefore lead to high compression; only $K(p_{w})$ grows to accommodate additional tasks, analogous to how a programming library would be used in novel ways to write a new program. Since DNNs already tend to learn simple representations (Blier and Ollivier,, 2018), our definition suggests that ordinary training in certain multi-task settings (those that reuse certain task components) might be a simple method for learning compositional representations. Indeed, this has long been hypothesized and observed empirically (Driscoll et al.,, 2024; Johnston and Fusi,, 2023; Lachapelle et al.,, 2023; Vafidis et al., 2024a, ; Maziarka et al.,, 2022; Vafidis et al., 2024b, ), especially in the case of disentangled representation learning, and could be verified more formally using our definition of representational compositionality.
Appendix G Prequential coding